# KAN-ODEs: Kolmogorov-Arnold Network Ordinary Differential Equations for Learning Dynamical Systems and Hidden Physics

### v1

@@ Abstract
Kolmogorov-Arnold Networks (KANs) as an alternative to Multi-layer perceptrons (MLPs) are a recent development demonstrating strong potential for data-driven modeling. This work applies KANs as the backbone of a Neural Ordinary Differential Equation framework, generalizing their use to the time-dependent and grid-sensitive cases often seen in scientific machine learning applications. The proposed KAN-ODEs retain the flexible dynamical system modeling framework of Neural ODEs while leveraging the many benefits of KANs, including faster neural scaling, stronger interpretability, and lower parameter counts when compared against MLPs. We demonstrate these benefits in three test cases: the Lotka-Volterra predator-prey model, Burgers’ equation, and the Fisher-KPP PDE. We showcase the strong performance of parameter-lean KAN-ODE systems generally in reconstructing entire dynamical systems, and also in targeted applications to the inference of a source term in an otherwise known flow field. We additionally demonstrate the interpretability of KAN-ODEs via activation function visualization and symbolic regression of trained results. The successful training of KAN-ODEs and their improved performance when compared to traditional Neural ODEs implies significant potential in leveraging this novel network architecture in myriad scientific machine learning applications.
@@ I Introduction
Dynamical system modeling is a key part of many branches of engineering and science. The development of data-driven approaches to infer models and laws that govern these systems is a major task in many fields and has given rise to numerous machine learning and neural network-based techniques. These techniques often have a tradeoff between the amount of prior knowledge embedded in the model versus the interpretability of the model: purely data-driven approaches tend to lack interpretability, and conversely interpretable methods typically require substantial knowledge or enforcement of preexisting governing laws. Novel methods that bridge this gap have the potential to not only infer high-accuracy models from data, but also provide useful insight into the dynamics of what are often unknown physical processes.
@@
Neural Ordinary Differential Equations (Neural ODEs) based on multilayer perceptrons (MLPs) [1, 2, 3] couple black box neural networks used as gradient getters with standard ODE solvers in order to learn the relationship between the current state of a system and its gradient. Originally proposed as an augmentation to the effectively fixed step sizes in residual neural networks used for general-purpose deep learning, Neural ODEs have also seen substantial use in engineering and scientific modeling thanks to their natural inclination for pairing with dynamical system solvers. By mapping the current state onto its gradients, Neural ODEs are inherently grid and timescale flexible, allowing for substantial interpolation and generalization of learned models compared with fixed-grid regression. Neural ODEs do not require prior knowledge of a system in order to infer dynamical models from large datasets thanks to their use of black box MLPs, though on the flip side of this benefit they typically deliver models with thousands to millions of parameters, making any deeper interpretation or extraction of physical insights difficult.
@@
Physics-Informed Neural Networks (PINNs) [4, 5, 6] are another key technique often used for machine learning based modeling of dynamical systems. This approach directly leverages the governing laws of a given system in the loss function of a neural network, allowing for training that not only aims to fit the training data but also aims to fit the known governing equations without having to directly solve what are typically expensive PDEs. These networks are significantly interpretable thanks to the governing equations used as a soft constraint in the training cycle, giving physical meaning to trends and patterns that emerge from the training process. They additionally typically require less training data than alternative black-box approaches, thanks to the significant amount of knowledge imparted by the governing equations. A recently developed approach with a similar incorporation of known physics is the Chemical Reaction Neural Network (CRNN) [7], which combines the efficient parameter optimization of Neural ODEs with a hard constraint on the system’s adherence to known physical laws. This allows the CRNN to move away from black-box style approaches and instead directly learn the system’s governing equation parameters from real [8, 9] and noisy [10, 11] experimental data. A downside of these physics informed and physics enforced techniques is their requirement of strong prior knowledge of the system’s governing laws. In cases where such knowledge is not available, other methods are required.
@@
The SINDy approach (Sparse Identification of Nonlinear Dynamics) [12, 13, 14, 15] tackles dynamical system modeling via sparse regression. In terms of prior knowledge and interpretability it appears to operate between Neural ODEs and physics-driven methods. Given training data and a set of candidate functions (polynomials, trigononometric functions, constant functions, etc.), it carries out sparse regression to parsimoniously select only the candidate functions needed to describe trends present in the data. The use of a presumed set of candidate functions gives SINDy models significant interpretability over pure black box methods like MLPS or Neural ODEs, while requiring less of the detailed physical knowledge behind the system dynamics as in physics-informed methods such as PINN and CRNN. This presumed set of candidate functions additionally often allows SINDy to train using significantly less data than traditional neural network approaches.
Kolmogorov-Arnold Networks (KANs) [16] are a recent development proposing an alternative to MLP-based networks. Instead of fixed activation functions with variable weights and biases, the activation functions themselves are made up of gridded basis functions, the scaling factors of which are trained in order to fit the data. Major selling points of KANs over MLPs are increased interpretability thanks to the direct learning of activation functions, faster convergence to low training loss with significantly fewer parameters thanks to their faster neural scaling laws, and the possibility of postprocessing via sparse regression to delivery SINDy-esque symbolic functions mapping the inputs to outputs. Unlike SINDy, however, these symbolic functions are not a presumed condition of training, and are instead an optional and modular postprocessing component, giving KANs greater flexibility and applicability to general use cases, which may benefit from black-box style models, symbolic function-based models, or even a combination of both. KANs have been extended to include time series predictions [17, 18] by directly mapping training data onto a future testing window. While these extensions allow for consideration of complex multivariate effects and concept drift, their rigid structure and direct treatment of fixed batches of data restricts their generalization outside of the explicit problem formulation under which they were trained.
@@
Here, we introduce the possibility of learning interpretable and modular ODE and PDE models for dynamical systems without prior knowledge or functional form assumptions by leveraging KANs as gradient-getters within the concept of Neural ODEs. In doing so, we propose a combined framework that benefits from the high-accuracy, parameter-learn, and strongly interpretable KAN network structure coupled with the general-purpose, grid-independent, and solver-flexible Neural ODE framework, A qualitative depiction of the comparison between the currently-proposed KAN-ODE method and the other discussed data-driven approaches for dynamical system modeling is shown in Fig. 1. While the current state of the art is generally situated on the diagonal of this plot, where increased prior knowledge fed into a machine learning method delivers better interpretability, we place KAN-ODEs slightly below the diagonal, thanks to their potential to deliver interpretable activations and even symbolic relationships without requiring prior knowledge or presumed functional forms in the training cycle. The ODE-based formulation is also inherently flexible, allowing for variable timesteps depending on the gradient landscape as well as application of the KAN-ODE as a component of a larger-scale model, for example as the source term in a nonlinear transport PDE.
@@ II.1 Kolmogorov Arnold Networks as Gradient Evaluators
Here, $\phi_{q,p}:~{}[0,1]\rightarrow\mathbb{R}$ are univariate functions, while $\Phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ composes these univariate functions to reconstruct $f(\textbf{x})$. The implementation in this work defines $\Phi_{q}$ as trainable weights to sum the univariate basis functions. Once these basis functions are chosen, Eq. 1 exactly describes a 2-layer KAN with hidden layer width of $12n+$ 1 (the upper bound of the outer summation in Eq. 1) and an output dimension of 1. The KAT is extended to a general KAN structure with $L$ layers of arbitrary width by exploiting the hierarchical structure of Eq. 1 [16].
@@
${\Phi}_{l}=\begin{pmatrix}\phi_{l,1,1}(\cdot)&\phi_{l,1,2}(\cdot)&\cdots&\phi_{l,1,n_{l}}(\cdot)\\ \phi_{l,2,1}(\cdot)&\phi_{l,2,2}(\cdot)&\cdots&\phi_{l,2,n_{l}}(\cdot)\\ \vdots&\vdots&&\vdots\\ \phi_{l,n_{l+1},1}(\cdot)&\phi_{l,n_{l+1},2}(\cdot)&\cdots&\phi_{l,n_{l+1},n_{l}}(\cdot)\\ \end{pmatrix}$ (3)
where $\phi$ is the activation function, $l$ is the index of a layer, $n_{l}$ and $n_{l+1}$ denote the numbers of nodes in the $l^{\text{th}}$ and $(l+1)^{\text{th}}$ layers, respectively. While Liu et al. [16] proposed a B-spline for this basis function, this study adopts the more computationally efficient Gaussian radial basis functions $\psi$ as proposed by Li [21], such that
$\phi_{l,\alpha,\beta}\left(\text{x}\right)$ $=\sum_{i=1}^{N}w^{r}_{l,\alpha,\beta,i}\cdot\psi\left(\lvert\lvert\text{x}-c_{i}\rvert\rvert\right)+w^{b}_{l,\alpha,\beta}\cdot b\left(\text{x}\right),$ (4)
@@
where $N$ is the number of gridpoints, $w^{r}_{l,\alpha,\beta,i}$ and $w^{b}_{l,\alpha,\beta}$ are the trainable weights for the RBF function $\psi(r)$ and the base activation function $b(\text{x})$ respectively (superscripts indicating which function the weights apply to), the subscripts $[\alpha,\beta]$ indicate the index of the matrix in Eq. 3 within a KAN layer $l$ that these weights apply to, $c_{i}$ is the center of the $i$-th RBF function, $r$ is the distance of x from the center, and $h$ is a spreading parameter defined as the gridpoint spacing. The inputs to each layer are normalized to be on the [-1, 1] gridded RBF domain as in [22], which avoids the costly alternative technique of [21] that requires periodically re-gridding the RBF networks in each layer to match the input domain. In addition to fully gridded RBF basis functions, the inputs and outputs of each layer are also directly connected with Swish residual activation functions such that $b(\text{x})=\text{x}\cdot\text{sigmoid}(\text{x})$ [23].
This discussion highlights the two input pathways for each layer: the RBF function on a grid of size N, and the direct (non-gridded) residual activations. The total number of parameters in the fully connected RBF pathway for a single layer is equal to the product of the input dimension, grid size, and output dimension (following from $w^{r}_{l,\alpha,\beta,i}$). The residual activation pathway has parameter count equal to just the product of the input and output dimensions (following from $w^{b}_{l,\alpha,\beta}$). For example, in a 2-layer KAN with a 2D input, 1D output, 10D hidden layer, and 5-point grid, the total number of parameters is [(2$\times$5$\times$10+2$\times$10)+(10$\times$5$\times$1+10$\times$1)]=180, when accounting for one $2\rightarrow 10$ layer and one $10\rightarrow 1$ layer in the two parenthetical terms, respectively. Expanded mathematical formulations for deep KANs and KANs with multivariate outputs become increasingly bulky and cumbersome, and are available in the extensive [16]. For convenience, we use the notation [$n_{l},n_{l+1},N$] to represent an $l$-th KAN layer for the remainder of this work, where $n_{l}$ is the input dimension, $n_{l+1}$ is the output dimension, and $N$ is the grid size. For example, a 2-input and 1-output KAN with a 10-node hidden layer and a 5-point grid is represented as a [2, 10, 5], [10, 1, 5].
@@ II.2 KAN-Ordinary Differential Equations
$\frac{d\mathbf{u}}{dt}=\mathbf{g}\left(\mathbf{u},t\right).$ (6)
@@
where $\bm{theta}$ is the collection of all trainable weights in the KAN. Here, the input and output dimensions of the KAN are fixed as the dimension of the state variables $\mathbf{u}$, and it is used as a gradient getter. Eq. 7 can be solved using a traditional ODE solver from an initial time $t_{0}$ to a desired time $t$ such that
@@
$\mathcal{L}\left(\bm{\theta}\right)$ $=\text{MSE}\left(\mathbf{u}^{\text{KAN}}\left(t,\bm{\theta}\right),\mathbf{u}^{\text{obs}}\left(t\right)\right)$ (9)
$=\frac{1}{N}\sum_{i=1}^{N}{\lvert\lvert\mathbf{u}^{\text{KAN}}\left(t_{i},\bm{\theta}\right)-\mathbf{u}^{\text{obs}}\left(t_{i}\right)\rvert\rvert^{2}},$
where $\mathbf{u}^{\text{KAN}}$ is the prediction, $\mathbf{u}^{\text{obs}}$ is the observation, and $N$ is the total number of time steps. To optimize $\mathcal{L}$, we require gradients with respect to $\bm{\theta}$. Generally, two options are available with the forward sensitivity method and the adjoint sensitivity method. In this paper, the adjoint sensitivity method is adopted to efficiently handle a potentially large Kolmogorov-Arnold network [24, 2, 1] thanks to its more advantageous scaling with model size. By introducing an adjoint state variable $\bm{\omega}$, the augmented system equation can be derived, which allows for gradient backpropagation through the ODE integrators.
$\frac{d}{dt}\left[\begin{array}[]{c}\mathbf{z}\\ \bm{\omega}\\ \frac{\partial\mathcal{L}}{\partial\bm{\theta}}\end{array}\right]=-\left[\begin{array}[]{ccc}1&\bm{\omega}^{T}&\bm{\omega}^{T}\end{array}\right]\left[\begin{array}[]{ccc}g&\frac{\partial g}{\partial\mathbf{z}}&\frac{\partial g}{\partial\bm{\theta}}\\ 0&0&0\\ 0&0&0\end{array}\right],$ (10)
where $\mathbf{z}(t)=\mathbf{u}(-t)$, $g=\text{KAN}$, and $T$ denotes the transpose. After Eq. 10 is computed, $\bm{\theta}$ is updated using a gradient descent method. A schematic depicting the overall training cycle is provided in Fig. 2.
@@
Figure 2: Schematic depicting the overall training cycle of a KAN-ODE. The inner loop in green leverages a KAN as a temporal gradient getter for the state vector to solve the ODE forward. Once a solution is generated, the outer blue loop computes the gradient of the loss function via the adjoint method to update the KAN activation functions.

### v2

@@ Abstract
Kolmogorov-Arnold networks (KANs) as an alternative to multi-layer perceptrons (MLPs) are a recent development demonstrating strong potential for data-driven modeling. This work applies KANs as the backbone of a neural ordinary differential equation (ODE) framework, generalizing their use to the time-dependent and temporal grid-sensitive cases often seen in dynamical systems and scientific machine learning applications. The proposed KAN-ODEs retain the flexible dynamical system modeling framework of Neural ODEs while leveraging the many benefits of KANs compared to MLPs, including higher accuracy and faster neural scaling, stronger interpretability and generalizability, and lower parameter counts. First, we quantitatively demonstrated these improvements in a comprehensive study of the classical Lotka-Volterra predator-prey model. We then showcased the KAN-ODE framework’s ability to learn symbolic source terms and complete solution profiles in higher-complexity and data-lean scenarios including wave propagation and shock formation, the complex Schrödinger equation, and the Allen-Cahn phase separation equation. The successful training of KAN-ODEs, and their improved performance compared to traditional Neural ODEs, implies significant potential in leveraging this novel network architecture in myriad scientific machine learning applications for discovering hidden physics and predicting dynamic evolution.
@@ I Introduction
Dynamical system modeling is a key part of many branches of engineering and science. Traditionally, governing equations (often partial differential equations (PDEs)) have been derived through extensive observations and careful selection of model equations. These processes, heavily relying on expert knowledge, have gradually been augmented and sometimes replaced with machine learning techniques by leveraging the data-fitting power in neural networks [1, 2, 3, 4]. Despite the potential to reduce the need for detailed expert knowledge, these data-driven modeling approaches are often subject to tradeoffs between the amount of prior knowledge embedded in the model versus the interpretability of the model: purely data-driven approaches tend to lack interpretability, and conversely, interpretable methods typically require substantial knowledge or enforcement of preexisting governing laws [1]. To mitigate this issue, there have been extensive efforts to develop novel modeling schemes that blend data-driven modeling with traditional dynamical system tools, including sparse identification of nonlinear dynamics (SINDy) [5], physics-informed neural network (PINNs) [6], and neural ordinary differential equations (Neural ODEs) [7]. Such studies that bridge this gap have the potential to not only infer high-accuracy models from data but also provide useful insight into the dynamics of what are often unknown physical processes.
@@
Neural ODEs have been developed to learn a continuous evolution of states [7, 8, 9], as opposed to the discrete sequences learned by residual networks [10, 11]. Chen et al. coupled black box neural networks used as gradient getters with standard ODE solvers in order to learn the relationship between the current state of a system and its gradient [7]. This mapping gives Neural ODEs inherent grid and timescale flexibility, allowing for substantial interpolation and generalization of learned models compared with typical fixed grid neural network approaches. Neural ODEs do not require prior knowledge of a system in order to infer dynamical models from large datasets thanks to their use of black box multi-layer perceptrons (MLPs). However, on the flip side of this benefit, they typically deliver models with thousands to millions of parameters, making any deeper interpretation or extraction of physical insights difficult.
@@
In contrast to Neural ODEs, physics-informed neural networks (PINNs) leverage the governing equations of the dynamical system, wherever available, in the loss function to train a traditional neural network that maps the time and spatial coordinates directly to the state variables [6, 12, 13]. This physics-embedded loss function allows for training that aims to fit both the data and the known governing equations without having to directly solve what are typically expensive PDEs. These networks are significantly interpretable thanks to the governing equations used as a soft constraint in the training cycle, giving physical meaning to trends and patterns that emerge from the training process. The inclusion of the governing equations as a soft constraint also allows the PINN to train with relatively small datasets, which is a useful benefit in many scientific and engineering applications with high costs of data acquisition [6]. Other frameworks have also been developed using a physics-based structure similar to PINN. The auto-regressive dense encoder-decoder model [14] encodes physics into a time-integration network scheme, while the physics-informed convolutional-recurrent network [15] runs PDE solutions through convolutional filters with physical penalty terms and hard-encoded boundary and initial conditions to efficiently develop models. Other works have also investigated the use of hard constraints for PINN-type models for initial and boundary conditions with strong results [16, 17]. PINN-type loss functions have also been included in Neural ODE frameworks [18]. Further combinations also exist in the realm of physics-based Neural ODEs, such as Chemical Reaction Neural Networks [19], which hard constrain Neural ODE frameworks to directly learn governing equation parameters from real [20, 21] and noisy [22, 23] experimental data. A downside of these physics-based techniques is their requirement of strong prior knowledge of the system’s governing laws. In cases where such knowledge is not available, other methods are required.
@@
The SINDy approach [5, 24, 25, 4] tackles dynamical system modeling via sparse regression. Given training data and a set of candidate functions (polynomials, trigononometric functions, constant functions, etc.), it carries out sparse regression to parsimoniously select only the candidate functions needed to describe trends present in the data. The use of a presumed set of candidate functions gives SINDy models significant interpretability over pure black box methods like MLPs or Neural ODEs. It also requires less of the detailed physical knowledge behind the system dynamics as in PINNs and other physics-based methods, and less data with a reduced model search space compared to Neural ODEs.
While most advances in data-driven modeling lie in MLPs and polynomial basis approaches, Liu et al. recently proposed Kolmogorov-Arnold networks (KANs) as an alternative to MLP-based networks [26]. Instead of learning weights and biases given fixed activation functions as in MLPs, the activation functions themselves are learnable in KANs with gridded basis functions and associated trainable scaling factors. Major advantages of KANs over MLPs are increased interpretability thanks to the direct learning of activation functions, faster convergence to low training loss with significantly fewer parameters thanks to their faster neural scaling laws, and the possibility of postprocessing via sparse regression to deliver human-readable expressions. With this potential, recent efforts have been dedicated to extending KANs for time-series predictions based on residual networks by directly mapping training data onto future testing windows [27, 28]. However, such a discrete time-step approach has a limited ability to predict the states where various time scales and data collection windows are involved. This research gap naturally motivates our work to embed KANs into Neural ODE-type dynamical system solvers, allowing us to resolve these drawbacks while preserving the interpretability and inference potential of KANs.
@@
In this study, we introduce the possibility of learning interpretable and modular ODE and PDE models for dynamical systems without prior knowledge or functional form assumptions by leveraging KANs as gradient-getters within the concept of Neural ODEs. In doing so, we propose a combined framework that benefits from the high-accuracy, parameter-learn, and strongly interpretable KAN network structure coupled with the general-purpose, grid-independent, and solver-flexible Neural ODE framework. A qualitative depiction of the comparison between the currently proposed KAN-ODE method and the other discussed data-driven approaches for dynamical system modeling is shown in Fig. 1. While the current state of the art is generally situated on the diagonal of this plot, where increased prior knowledge fed into a machine learning method delivers better interpretability, we place KAN-ODEs slightly below the diagonal, thanks to their potential to deliver interpretable activations and even symbolic relationships without requiring prior knowledge or presumed functional forms in the training cycle. The ODE-based formulation is also inherently flexible, allowing for variable timesteps depending on the gradient landscape as well as the application of the KAN-ODE as a component of a larger-scale model, for example as the source term in a nonlinear transport PDE. We demonstrated the promise of KAN-ODEs through a comprehensive comparison against standard Neural ODEs, as well as an array of tests at different scales and levels of complexity including a predator-prey ODE, wave propagation and shock formation PDEs, the complex Schrödinger equation, and a phase separation PDE. We found that KAN-ODEs succeed in all of these cases and are able to reconstruct complete temporal profiles even when given sparse training data. KAN-ODEs with a small number of parameters were able to converge with remarkable accuracy, and sparsity or symbolic relationships imposed during the training cycle or during postprocessing were able to give KAN-ODEs substantial generalization capability. Overall, we found that this encoding of KANs into the Neural ODE framework is a promising step toward small-data and low-knowledge yet interpretable machine learning for dynamical system modeling.
@@ II.1 Kolmogorov Arnold Networks as Gradient Evaluators
Here, $\phi_{q,p}:~{}[0,1]\rightarrow\mathbb{R}$ are univariate functions, while $\Phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ composes these univariate functions to reconstruct $f(\textbf{x})$. The implementation in this work defines $\Phi_{q}$ as trainable weights to sum the univariate basis functions. Once these basis functions are chosen, Eq. 1 exactly describes a 2-layer KAN with a hidden layer width of $12n+$ 1 (the upper bound of the outer summation in Eq. 1) and an output dimension of 1. The KAT is extended to a general KAN structure with $L$ layers of arbitrary width by exploiting the hierarchical structure of Eq. 1 [26].
@@
${\Phi}_{l}=\begin{pmatrix}\phi_{l,1,1}(\cdot)&\phi_{l,1,2}(\cdot)&\cdots&\phi_{l,1,n_{l}}(\cdot)\\ \phi_{l,2,1}(\cdot)&\phi_{l,2,2}(\cdot)&\cdots&\phi_{l,2,n_{l}}(\cdot)\\ \vdots&\vdots&&\vdots\\ \phi_{l,n_{l+1},1}(\cdot)&\phi_{l,n_{l+1},2}(\cdot)&\cdots&\phi_{l,n_{l+1},n_{l}}(\cdot)\\ \end{pmatrix},$ (3)
where $\phi$ is the activation function, $l$ is the index of a layer, and $n_{l}$ and $n_{l+1}$ denote the numbers of nodes in the $l^{\text{th}}$ and $(l+1)^{\text{th}}$ layers, respectively. While Liu et al. [26] originally proposed a B-spline for the gridded basis functions that comprise the various activations $\phi$, this study adopts the more computationally efficient Gaussian radial basis functions (RBFs) $\psi$ as proposed by Li [31], such that
$\phi_{l,\alpha,\beta}\left(\text{x}\right)$ $=\sum_{i=1}^{N}w^{\psi}_{l,\alpha,\beta,i}\cdot\psi\left(\lvert\lvert\text{x}-c_{i}\rvert\rvert\right)+w^{b}_{l,\alpha,\beta}\cdot b\left(\text{x}\right),$ (4)
@@
where $N$ is the number of gridpoints, $w^{\psi}_{l,\alpha,\beta,i}$ and $w^{b}_{l,\alpha,\beta}$ are the trainable weights for the RBF function $\psi(r)$ and the base activation function $b(\text{x})$ respectively (superscripts indicating which function the weights apply to), the subscripts $[\alpha,\beta]$ indicate the index of the matrix in Eq. 3 within a KAN layer $l$ that these weights apply to, $c_{i}$ is the center of the $i$-th RBF function, $r$ is the distance of x from the center, and $h$ is a spreading parameter defined as the gridpoint spacing. The inputs to each layer are normalized to be on the [-1, 1] gridded RBF domain as in [32], which avoids the costly alternative technique of [26] that requires periodically re-gridding the RBF networks in each layer to match the input domain. In addition to fully gridded RBF basis functions, the inputs and outputs of each layer are also directly connected with Swish residual activation functions such that $b(\text{x})=\text{x}\cdot\text{sigmoid}(\text{x})$ [33].
Equation 4 therefore defines two input pathways for each layer: the gridded RBF pathway (first term), and the direct (non-gridded) residual activation pathway (second term). The RBF function connects each input node with each output node using a number of basis functions equal to the size of the grid. The number of parameters required for the RBF pathway of a single layer is therefore equal to the product of the input dimension, output dimension, and grid size (following as well from $w^{\psi}_{l,\alpha,\beta,i}$, where for a single layer $w^{\psi}_{l}$ the remaining three indices cover the input, output, and grid indices). Similarly, the residual activation pathway, which directly connects inputs to outputs without a grid, simply requires parameters equal to the product of the input and output dimensions (likewise see $w^{b}_{l,\alpha,\beta}$, which includes the input and output indices $\alpha$ and $\beta$ but not the grid index $i$). For example, in a 2-layer KAN with a 2D input, 1D output, 10D hidden layer, and 5-point grid, the numbers of parameters are 2$\times$5$\times$10+2$\times$10 for the input $\rightarrow$ hidden layer, and 10$\times$5$\times$1+10$\times$1 for the hidden $\rightarrow$ output layer, respectively, resulting in 180 parameters in total.
Expanded mathematical formulations for KANs with multivariate outputs become increasingly bulky and cumbersome and are available in the extensive [26]. For convenience, we use the notation [$n_{l},n_{l+1},N$] to represent an $l$-th KAN layer for the remainder of this work, where $n_{l}$ is the input dimension, $n_{l+1}$ is the output dimension, and $N$ is the grid size. For example, a 2-input and 1-output KAN with a 10-node hidden layer and a 5-point grid is represented as ”[2, 10, 5], [10, 1, 5]”, for the connections between the input and hidden layer, and the hidden and output layer, respectively.
@@ II.2 KAN-Ordinary Differential Equations
$\frac{d\mathbf{u}}{dt}=\mathbf{g}\left(\mathbf{u},t\right),$ (6)
@@
where $\bm{\theta}$ is the collection of all trainable weights in the KAN. Here, the input and output dimensions of the KAN are fixed as the dimension of the state variables $\mathbf{u}$, and it is used as a gradient getter. Equation  7 can be solved using a traditional ODE solver from an initial time $t_{0}$ to a desired time $t$ such that
@@
$\mathcal{L}\left(\bm{\theta}\right)=\text{MSE}\left(\mathbf{u}^{\text{KAN}}\left(t,\bm{\theta}\right),\mathbf{u}^{\text{obs}}\left(t\right)\right)=\frac{1}{N}\sum_{i=1}^{N}{\lvert\lvert\mathbf{u}^{\text{KAN}}\left(t_{i},\bm{\theta}\right)-\mathbf{u}^{\text{obs}}\left(t_{i}\right)\rvert\rvert^{2}},$ (9)
where $\mathbf{u}^{\text{KAN}}$ is the prediction, $\mathbf{u}^{\text{obs}}$ is the observation, and $N$ is the total number of time steps. To optimize $\mathcal{L}$, we require gradients with respect to $\bm{\theta}$. Generally, two options are available with the forward sensitivity method [8, 34] and the adjoint sensitivity method [7, 35]. In this paper, the adjoint sensitivity method is adopted to efficiently handle a potentially large Kolmogorov-Arnold network [35, 8, 7] thanks to its more advantageous scaling with model size. By introducing an adjoint state variable $\bm{\omega}$, the augmented system equation can be derived, which allows for gradient backpropagation through the ODE integrators.
$\frac{d}{dt}\left[\begin{array}[]{c}\mathbf{z}\\ \bm{\omega}\\ \frac{\partial\mathcal{L}}{\partial\bm{\theta}}\end{array}\right]=-\left[\begin{array}[]{ccc}1&\bm{\omega}^{T}&\bm{\omega}^{T}\end{array}\right]\left[\begin{array}[]{ccc}\mathbf{g}&\frac{\partial\mathbf{g}}{\partial\mathbf{z}}&\frac{\partial\mathbf{g}}{\partial\bm{\theta}}\\ 0&0&0\\ 0&0&0\end{array}\right],$ (10)
where $\mathbf{z}(t)=\mathbf{u}(-t)$, $\mathbf{g}=\text{KAN}$, and $T$ denotes the transpose. After Eq. 10 is computed, $\bm{\theta}$ is updated using a gradient descent method. A schematic depicting the overall training cycle is provided in Fig. 2.
@@
Figure 2: Schematic depicting the overall training cycle of a KAN-ODE. The loop in green leverages a KAN as a temporal gradient getter for the state vector to solve the ODE forward. Once a solution is generated, the blue loop computes the gradient of the loss function via the adjoint method to update the KAN activation functions.